OpenStudy (anonymous):

solve log2+logx=log32

OpenStudy (anonymous):

use properties of logarithms\[\log a + \log b = \log ab\]we use this property for the left side\[\log 2 + \log x = \log 32\]\[\log 2x = \log 32\]\[2x = 32\]\[x = 16\]